21. If one root of the quadratic equation 2x² + kx – 6 = 0 is 2, the value of k is
(a) 1
(b) -1
(c) 2
(d) -2

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: b Explaination:Reason: Scice x = 2 is a root of the equation 2x² + kx -6 = 0 ∴ 2(2)² +k(2) – 6 = 0 ⇒ 8 + 2k – 6 = 0 ⇒ 2k = -2 ∴ k = -1